OpenStudy (jotopia34):

I don't understand this, please help. Determine if the series is convergent or divergent: If convergent, find the sum othe series. (sigma)from n=0 to infinity of (-1)^n*3/4^n

OpenStudy (anonymous):

This series \[\huge \sum_{n=0}^{\infty}(-1)^n\left(\frac{3}{4}\right)^n\]?

OpenStudy (anonymous):

converges like anything

OpenStudy (anonymous):

Except something divergent, @satellite73 :)

OpenStudy (anonymous):

true that

OpenStudy (jotopia34):

okay smarties,thanks. But how do I show this? What "test" do I use, I just don't understand it :(

OpenStudy (anonymous):

a) \((\frac{3}{4})^n\to 0\) as \(n\to \infty\) and the series alternates, so that is enough

OpenStudy (anonymous):

Who needs tests? Without the (-1)^n, it's just a geometric series with a common ratio less than 1 So it's absolutely convergent :D

OpenStudy (jotopia34):

but it has the (-1)^n, so why isnt that significant?

OpenStudy (anonymous):

b)\( \sum(\frac{3}{4})^n=\frac{1}{1-\frac{3}{4}}=4\)

OpenStudy (jotopia34):

well peter pan, "Professor Wendy" requires tests

OpenStudy (anonymous):

I mean, the series is absolutely convergent without the (-1)^n so with the (-1)^n, it would still be convergent :)

OpenStudy (anonymous):

it makes it smaller so it converges for sure

OpenStudy (anonymous):

That's annoying of her :) Use the Alternating Series test, see if the sequence converges to zero :)

OpenStudy (anonymous):

or compare with \[\sum(\frac{3}{4})^n=\frac{1}{1-\frac{3}{4}}=4\]

OpenStudy (anonymous):

since it is smaller than 4, it converges

OpenStudy (jotopia34):

\[\sum_{n=0}^{\infty}(-1)^n(\frac{ 3 }{ 4^n })\]

OpenStudy (anonymous):

Can only compare positive series with the comparison test, @satellite73 Naughty naughty...

OpenStudy (anonymous):

For instance, this series can also be compared with your series @satellite73 but it's obviously divergent \[\huge \sum_{n=1}^{\infty}-\frac1n\]

OpenStudy (anonymous):

If you have a series \[\huge \sum_{n=k}^{\infty}(-1)^na_n\]

OpenStudy (jotopia34):

I understand the p-series. But why would I ignore the (-1)^n part?

OpenStudy (anonymous):

Where a_n is a positive series, if the limit of a_n as n becomes really really big is zero, the series \[\huge \sum_{n=k}^{\infty}(-1)^na_n\]

OpenStudy (anonymous):

is convergent :)

OpenStudy (anonymous):

But if you don't like that, there's another way. :)

OpenStudy (jotopia34):

Thats super. TY. Then, what is the sum? Is that the limit?

OpenStudy (jotopia34):

What is the other way?

OpenStudy (anonymous):

No. I'm afraid getting the sum isn't THAT simple :)

OpenStudy (jotopia34):

crap, ok how?

OpenStudy (anonymous):

The other way is the root test :) Given a series \[\huge \sum_{n=k}^{\infty}a_n\]if \[\huge \lim_{n\rightarrow \infty}\sqrt[n]{|a_n|}<1\]

OpenStudy (anonymous):

Then the series is absolutely convergent :)

OpenStudy (jotopia34):

yeah, no on the root test here

OpenStudy (anonymous):

Okay, then stick to the Alternating series test :)

OpenStudy (jotopia34):

what is the sum however? I am not sure

OpenStudy (anonymous):

why can't i compare?

OpenStudy (anonymous):

Actually, our series may be written as \[\huge \sum_{n=0}^{\infty}\left(-\frac{3}{4}\right)^n\]

OpenStudy (anonymous):

@satellite73 Were you using the direct Comparison test?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Well, you may only compare two series \[\huge \sum_{n=k}^{\infty}a_n\]and \[\huge \sum_{n=k}^{\infty}b_n\] If both a_n and b_n are positive series.

OpenStudy (anonymous):

oh i see, the negative terms. in any case the terms go to zero, not minus infinity or something

OpenStudy (anonymous):

I suppose, the absolute values of the terms of the series are exactly the same as the terms of a known convergent series, comparison test, you get equality, you get absolute convergence, Boom, convergence ^.^

OpenStudy (jotopia34):

The sum!!!

OpenStudy (anonymous):

add two series might work

OpenStudy (jotopia34):

Need the sum!!!

OpenStudy (anonymous):

I told you, @jotopia34 The series is just like \[\huge \sum_{n=0}^{\infty}\left(-\frac{3}{4}\right)^n\]

OpenStudy (anonymous):

oh better idea

OpenStudy (anonymous):

And it's a geometric series with first term 1 and common ratio -3/4

OpenStudy (jotopia34):

so I do a/1-r

OpenStudy (anonymous):

And while you're thinking about that, let's play with infinite geometric series :) The sum can be written \[\Large S=a+ar+ar^2+ar^3+ar^4+...\]\[\Large S-a=ar+ar^2+ar^3+ar^4+ar^5+...\]\[\Large S-a=r(a+ar+ar^2+ar^3+ar^4+...\]\[\Large S-a=rS\]\[\Large S-rS=a\]\[\Large S(1-r)=a\]\[\Large S=\frac{a}{1-r}\]

OpenStudy (anonymous):

Oh, sorry, what were we talking about? ^.^

OpenStudy (anonymous):

lol reinventing the wheel again

OpenStudy (anonymous):

Because I can >:)

OpenStudy (jotopia34):

lol, so I take it the answer is yes to my sum solving proposal

OpenStudy (anonymous):

hey, @satellite73 What do you mean by "again" ? >.>

OpenStudy (anonymous):

And yes, @jotopia34 ^.^